OpenStudy (anonymous):

solve each system by substitution Please show how you got it and also how you found Y... Thanks :) x+7y=0 2x-8y=22

OpenStudy (anonymous):

x=-7y (took the first equation and solved y for x) 2(-7y)-8y=22 (plug in the new equation for x) -14y-8y=22 (simplify) -22y=22 (simplify) y=-1(divide both sides by -22)

OpenStudy (anonymous):

x+7(-1)=0 (plug in -1 for y) x-7=0 x=7

OpenStudy (anonymous):

why did the 7 turn (-)

OpenStudy (anonymous):

I plugged in the new y value, which we found out was negative one. 7*-1=-7

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yup!